Assuming the following ratios are constant, what is the sustainable growth rate?

Profit margin = 7.3 percent; Capital intensity (total assets/sales) = 0.80; Debt-equity ratio = 0.95; Net income = $73,000; Dividends = $24,000.
Answer #1

ROE =Profit Margin* Capital Intensity * (1+Debt/Equity Ratio) =7.3%*0.80*(1+0.95) =11.388%

Retention Ratio =(1-Dividends/Net Income) =(1-24000/73000) =0.6712

Sustainable growth rate =ROE*Retention Ratio/(1-ROE*Retention ratio) =11.388%*0.6712/(1-11.388%*0.6712) =8.28%